OpenStudy (anonymous):

Find the exact value of Cos(A+B) if SinA=-4/5 and SinB=-12/13 where both A and B are in quadrant 4

OpenStudy (zzr0ck3r):

\(\cos(a+b)=\cos(a)\cos(b)+\sin(a)\sin(b)\)

OpenStudy (zzr0ck3r):

\(\cos^2(a)+\sin^2(a) = 1\) use this along with the given info to figure out cos(a) then do the same thing with b.